# Decompose and Compare Consistency: Measuring VLMs’ Answer Reliability via Task-Decomposition Consistency Comparison

### v1

@@ Abstract
Despite tremendous advancements, current state-of-the-art Vision-Language Models (VLMs) are still far from perfect. They tend to hallucinate and may generate biased responses. In such circumstances, having a way to assess the reliability of a given response generated by a VLM is quite useful. Existing methods, such as estimating uncertainty using answer likelihoods or prompt-based confidence generation, often suffer from overconfidence. Other methods use self-consistency comparison but are affected by confirmation biases. To alleviate these, we propose Decompose and Compare Consistency (DeCC) for reliability measurement. By comparing the consistency between the direct answer generated using the VLM’s internal reasoning process, and the indirect answers obtained by decomposing the question into sub-questions and reasoning over the sub-answers produced by the VLM, DeCC measures the reliability of VLM’s direct answer. Experiments across six vision-language tasks with three VLMs show DeCC’s reliability estimation achieves better correlation with task accuracy compared to the existing methods.
@@ 3.1 Task Decomposition
First, the decomposer, which could be any VLM, decomposes the question into a sequence of sub-questions conditioned on . The candidate VLM then answers these sub-questions, resulting in a sequence of sub-QA pairs. Next, the candidate VLM and a separate LLM, acting as two independent agents, reason over the sub-QA pairs and , yielding VLM’s reasoned answer and LLM’s reasoned answer . To enhance robustness, we also experiment with a two-iteration decomposition process. In the second iteration, sub-QA pairs from the first iteration, along with and , are used to guide the decomposer in generating additional sub-questions. The candidate VLM answers these new sub-questions, conditioned on and previous sub-QA pairs, resulting in new sub-QA pairs containing more information. Finally, both agents reason over all sub-QA pairs from both iterations to provide their updated reasoned answers, and .
@@ 4.2 Existing Methods Used for Comparison
Perplexity of Direct Answer: Calculate the mean perplexity over tokens of the direct answer and use a threshold to determine reliability. If perplexity exceeds the threshold, is 0 otherwise 1. Generated Numerical Confidence: Prompt the VLM to generate a confidence value along with the answer, formatted as ‘Answer: X. Confidence: X%’. A threshold determines reliability. Generated Linguistic Confidence: Prompt the VLM to state ‘I am confident/not confident in this answer.’ Self-Consistency based on Paraphrase: Prompt a VLM to paraphrase the original question into four variations. If or more paraphrased answers differ from the direct answer, is 0 otherwise 1. 111We select the best threshold and for each VLM based on the Brier Score (results in Tables 2 and 3).
4.3 Results
@@
We conduct experiments on six vision-language tasks222All datasets are multiple-choice (model generates the index of the choice) except for MMMU, whose answers are very short. We use string matching for consistency comparison., covering commonsense reasoning, fine-grained compositional reasoning, and science understanding (see Appendix A.1 for dataset descriptions). We evaluate three state-of-the-art VLMs: LLaVA1.5-7B Liu et al. (2023), Idefics2-8B Laurençon et al. (2024), and InternVL1.5-25.5B Chen et al. (2024b) (see Appendix A.2 for implementation details). The overall results are shown in Table 1. DeCC achieves the best and second-best mean performance (mean across datasets) on Brier Score and Effective Reliability. DeCC reduces the relative mean Brier Score by 8.7% on LLaVA, 14.3% on Idefics2, and 2.9% on InternVL compared to the best existing methods. DeCC also increases relative mean Effective Reliability by 16.5% on LLaVA, 25.6% on Idefics2, and 1.7% on InternVL. We observe that with increasing VLM size, the performance of most methods improves, suggesting that reliability measurement is correlated with VLMs’ capabilities. For the effectiveness of DeCC’s different consistency comparison settings, we observe an interesting trend: (1) For weaker VLMs, i.e., LLaVA, LLM Agent Consistency achieves the best performance, likely because VLMs struggle to reason over the sub-QA pairs and suffer from confirmation biases. (2) For stronger VLMs, i.e. Idefics2, Multi-Agent Consistency performs the best suggesting that the VLM and LLM reasoners complement each other. (3) For the strongest VLMs, i.e. InternVL, VLM Agent Consistency (self-consistency) achieves the best performance, as the VLM can effectively leverage the information contained in sub-QA pairs. Overall, the effectiveness of different consistency comparison settings correlates with the candidate VLM’s capabilities.
@@ A.1 Datasets
We conduct experiments on six vision-language tasks: SNLI-VE (Xie et al., 2019), VCR (Zellers et al., 2019), A-OKVQA (Schwenk et al., 2022), Winoground (Thrush et al., 2022), MMMU Yue et al. (2023), and MathVista Lu et al. (2024). SNLI-VE requires VLMs to identify whether the relationship between the given image premise and text hypothesis is entailment, neutral, or contradiction. Visual Commonsense Reasoning (VCR) requires higher-order cognition and commonsense reasoning of VLMs. It provides an image and a question about certain objects in the image, along with four candidate answers, where the VLMs need to choose the correct answer. We add rectangles of different colors to the image and indicate the corresponding object’s index in the upper right corner of each rectangle to distinguish the objects. A-OKVQA is an augmented successor of OK-VQA (Marino et al., 2019) and requires a broad base of commonsense and world knowledge to answer questions. Four candidate answers are provided along with each question. Winoground (Wino.) is proposed for measuring vision-linguistic compositional reasoning. It contains two images and two captions. The model needs to correctly match the captions to the images, but crucially, both captions contain an identical set of words, only in a different order. MMMU is designed to evaluate VLMs on massive multi-discipline tasks demanding college-level subject knowledge and deliberate reasoning. Several candidate answers are provided along with each question. MathVista focuses on mathematical reasoning in visual contexts. We treat all datasets except for MathVista as multiple-choice QA tasks. For evaluation:
@@ A.2 Implement Details
We use InternVL-1.5 Chen et al. (2024b) as the decomposer for decomposition and question paraphrasing. For decomposition, we employ few-shot prompting by randomly selecting four samples from SNLI-VE and ScienceQA, with manually written decomposition processes as guidance. The few-shot prompt for decomposition is provided in Table 4. Only text is used in the few-shot prompt, without images. The decomposer determines the number of sub-questions needed. The few-shot prompt for the second-iteration decomposition is shown in Table 5 For paraphrasing, we use the same samples with manually written paraphrased questions. The few-shot prompt for paraphrasing is provided in Table 6. The remaining datasets are approached with a zero-shot strategy. We use OpenHermes-2.5-Mistral-7B333https://huggingface.co/teknium/OpenHermes-2.5-Mistral-7B as the LLM for reasoning. We evaluate three VLMs: LLaVA1.5-7B Liu et al. (2023), Idefics2-8B Laurençon et al. (2024), and InternVL Chen et al. (2024b), all operating under a zero-shot setting across all datasets. Since all datasets are multiple-choice QA tasks or short answers, we use string matching for answer consistency. For baseline threshold settings:
@@ A.4 Case Study
Main Question: Is this statement entailment, neutral, or contradiction based on the image? Statement: ’A professor is late to class’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 1: Is this statement entailment, neutral, or contradiction based on the image? Statement: ’A teacher is late to class’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 2: Is this statement entailment, neutral, or contradiction based on the image? Statement: ’A professor is tardy to class’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 3: Is this statement entailment, neutral, or contradiction based on the image? Statement: ’A professor is not on time for class’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 4: Is this statement entailment, neutral, or contradiction based on the image? Statement: ’A teacher is not punctual for class’ Options: A: entailment, B: neutral, C: contradiction. |
@@
Main Question: Is this statement entailment, neutral, or contradiction based on the image? Statement: ’Two children play in the park.’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 1: Is this statement entailment, neutral, or contradiction based on the image? Statement: ’Two kids play in the park.’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 2: Is this statement entailment, neutral, or contradiction based on the image? Statement: ’Two children are playing in the park.’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 3: Is this statement entailment, neutral, or contradiction based on the image? Statement: ’Two kids are playing in the park.’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 4: Is this statement entailment, neutral, or contradiction based on the image? Statement: ’There are two children playing in the park.’ Options: A: entailment, B: neutral, C: contradiction. |

### v3

@@ Abstract
Despite tremendous advancements, current state-of-the-art Vision-Language Models (VLMs) are still far from perfect. They tend to hallucinate and may generate biased responses. In such circumstances, having a way to assess the reliability of a given response generated by a VLM is quite useful. Existing methods, such as estimating uncertainty using answer likelihoods or prompt-based confidence generation, often suffer from overconfidence. Other methods use self-consistency comparison but are affected by confirmation biases. To alleviate these, we propose Decompose and Compare Consistency (DeCC) for reliability measurement. By comparing the consistency between the direct answer generated using the VLM’s internal reasoning process, and the indirect answers obtained by decomposing the question into sub-questions and reasoning over the sub-answers produced by the VLM, DeCC measures the reliability of VLM’s direct answer. Experiments across six vision-language tasks with three VLMs show DeCC’s reliability estimation achieves better correlation with task accuracy compared to the existing methods. The code is publicly available at https://github.com/MyLittleChange/DeCC.
@@ 3.1 Task Decomposition
First, the decomposer, which could be any VLM, decomposes the question into a sequence of sub-questions conditioned on . The candidate VLM then answers these sub-questions, resulting in a sequence of sub-QA pairs. Next, the candidate VLM and a separate LLM, acting as two independent agents, reason over the sub-QA pairs and , yielding VLM’s reasoned answer and LLM’s reasoned answer . To enhance robustness, we also experiment with a two-iteration decomposition process. In the second iteration, sub-QA pairs from the first iteration, along with and , are used to guide the decomposer in generating additional sub-questions. The candidate VLM answers these new sub-questions, conditioned on and previous sub-QA pairs, resulting in new sub-QA pairs containing more information. Finally, both agents reason over all sub-QA pairs from both iterations to provide their updated reasoned answers, and 111Single quotation mark (′) annotates the first iteration and double quotation mark (′′) annotates the second iteration..
@@ 4.3 Existing Methods Used for Comparison
Perplexity of Direct Answer: Calculate the mean perplexity over tokens of the direct answer and use a threshold to determine reliability. If perplexity exceeds the threshold, is 0 otherwise 1. Generated Numerical Confidence: Prompt the VLM to generate a confidence value along with the answer, formatted as ‘Answer: X. Confidence: X%’. A threshold determines reliability. Generated Linguistic Confidence: Prompt the VLM to state ‘I am confident/not confident in this answer.’ Self-Consistency based on Paraphrase: Prompt a VLM to paraphrase the original question into four variations. If or more paraphrased answers differ from the direct answer, is 0 otherwise 1. 222We select the best threshold and for each VLM based on the Brier Score (results in Tables 4 and 5).
Method | VCR | AOKVQA | Winoground | Mean | ||||
BS | ER | BS | ER | BS | ER | BS | ER | |
Perplexity of Direct Answer | 34.1 | 28.2 | 19.9 | 63.2 | 29.8 | 43.5 | 27.9 | 45.0 |
Generated Numerical Confidence | 37.7 | 25.3 | 36.3 | 46.7 | 25.3 | 49.1 | 33.1 | 40.4 |
Generated Linguistic Confidence | 40.2 | 22.1 | 25.2 | 56.6 | 26.8 | 45.6 | 30.7 | 41.4 |
Self-Consistency based on Paraphrase | 31.6 | 30.4 | 16.3 | 66.5 | 28.9 | 43.8 | 25.6 | 46.9 |
[…]
VLM Agent Consistency | 30.5 | 31.6 | 13.9 | 69.2 | 22.6 | 50.4 | 22.3 | 50.4 |
LLM Agent Consistency | 37.9 | 24.4 | 26.3 | 56.5 | 35.3 | 38.0 | 33.2 | 39.6 |
MultiAgent Consistency | 33.0 | 27.9 | 19.6 | 65.5 | 29.5 | 44.1 | 27.4 | 45.8 |
[…]
VLM Agent Consistency | 31.2 | 31.1 | 13.2 | 69.9 | 22.9 | 50.3 | 22.4 | 50.5 |
LLM Agent Consistency | 38.1 | 24.2 | 24.6 | 58.5 | 35.1 | 38.2 | 32.6 | 40.3 |
MultiAgent Consistency | 32.7 | 29.6 | 18.5 | 64.6 | 27.8 | 47.2 | 26.3 | 47.1 |
[…]
VLM Agent Consistency | 31.7 | 30.6 | 14.1 | 69.0 | 23.3 | 49.9 | 23.0 | 49.9 |
LLM Agent Consistency | 38.6 | 23.7 | 26.8 | 56.3 | 33.8 | 39.5 | 33.0 | 39.9 |
MultiAgent Consistency | 33.7 | 28.6 | 20.9 | 63.3 | 28.4 | 44.9 | 27.7 | 45.6 |
Benchmark | Questions | Question Types |
per Sample | per Sample | |
SNLI | 4.1 | 2.1 |
VCR | 5.2 | 2.2 |
AOKVQA | 4.4 | 2.1 |
Winoground | 4.3 | 2.5 |
MMMU | 5.9 | 2.4 |
MathVista | 4.4 | 2.3 |
4.4 Main Results
@@
We conduct experiments on six vision-language tasks333All datasets are multiple-choice except for MMMU and MathVista, whose answers are very short. We use string matching for consistency comparison., covering commonsense reasoning, fine-grained compositional reasoning, and science understanding (see Appendix A.1 for dataset descriptions). We evaluate three state-of-the-art VLMs: LLaVA1.5-7B Liu et al. (2023), Idefics2-8B Laurençon et al. (2024), and InternVL1.5-25.5B Chen et al. (2024b) (see Appendix A.2 for implementation details). The overall results are shown in Table 1. DeCC achieves the best and second-best mean performance (mean across datasets) on Brier Score and Effective Reliability. DeCC reduces the relative mean Brier Score by 8.7% on LLaVA, 14.3% on Idefics2, and 2.9% on InternVL compared to the best existing methods. DeCC also increases relative mean Effective Reliability by 16.5% on LLaVA, 25.6% on Idefics2, and 1.7% on InternVL. We observe that with increasing VLM size, the performance of most methods improves, suggesting that reliability measurement is correlated with VLMs’ capabilities. For the effectiveness of DeCC’s different consistency comparison settings, we observe an interesting trend: (1) For weaker VLMs, i.e., LLaVA, LLM Agent Consistency achieves the best performance, likely because VLMs struggle to reason over the sub-QA pairs and suffer from confirmation biases. (2) For stronger VLMs, i.e. Idefics2, Multi-Agent Consistency performs the best suggesting that the VLM and LLM reasoners complement each other. (3) For the strongest VLMs, i.e. InternVL, VLM Agent Consistency (self-consistency) achieves the best performance, as the VLM can effectively leverage the information contained in sub-QA pairs. Overall, the effectiveness of different consistency comparison settings correlates with the candidate VLM’s capabilities.
4.5 Further Analysis
Decoding Strategy
The decoding strategy of decomposer can influence the generated sub-questions, which in turn impacts DeCC’s overall performance. We conducted experiments using sampling-based decoding with two different temperature settings, and , while keeping the nucleus sampling Holtzman et al. (2020) probability fixed at for both. The results are presented in Table 2, using Idefics2-8B as the candidate VLM. Across all decoding strategies and temperature settings, DeCC consistently outperforms the baselines, demonstrating its effectiveness.
Question Type Analysis
Previous studies have shown that VLMs can exhibit biases toward certain question types. For instance, they respond “yes” to over 80% of queries about non-existent objects, such as in prompts like “Is there an object in this image?” Wang et al. (2023). If the decomposed sub-questions retain the same type as the original question, they may inherit the same bias, affecting evaluation reliability. To investigate this, we conduct the question type analysis to assess whether DeCC is prone to this issue 444We categorize questions into types using string matching, with categories including: yes/no, color, number, how, why, what/which, when, where, who, and others.. Table 3 shows the number of questions and question types per sample across benchmarks. DeCC generates at least 2.1 distinct question types per sample for all benchmarks, reducing the impact of VLM biases.
Additional Analysis
@@ 5 Conclusion
Acknowledgement
We are grateful to Mila’s IDT team for their technical support with the computational infrastructure. We thank Rabiul Awal for his constructive feedback. During this project, Aishwarya Agrawal was supported by the Canada CIFAR AI Chair award.
@@ A.1 Datasets
SNLI-VE requires VLMs to identify whether the relationship between the given image premise and text hypothesis is entailment, neutral, or contradiction. Visual Commonsense Reasoning (VCR) requires higher-order cognition and commonsense reasoning of VLMs. It provides an image and a question about certain objects in the image, along with four candidate answers, where the VLMs need to choose the correct answer. We add rectangles of different colors to the image and indicate the corresponding object’s index in the upper right corner of each rectangle to distinguish the objects. A-OKVQA is an augmented successor of OK-VQA (Marino et al., 2019) and requires a broad base of commonsense and world knowledge to answer questions. Four candidate answers are provided along with each question. Winoground (Wino.) is proposed for measuring vision-linguistic compositional reasoning. It contains two images and two captions. The model needs to correctly match the captions to the images, but crucially, both captions contain an identical set of words, only in a different order. MMMU is designed to evaluate VLMs on massive multi-discipline tasks demanding college-level subject knowledge and deliberate reasoning. Several candidate answers are provided along with each question. MathVista focuses on mathematical reasoning in visual contexts. We treat all datasets except for MMMU and MathVista as multiple-choice QA tasks. For evaluation:
@@ A.2 Implement Details
We use InternVL-1.5 Chen et al. (2024b) as the decomposer for decomposition and question paraphrasing. For decomposition, we employ few-shot prompting by randomly selecting four samples from SNLI-VE and ScienceQA, with manually written decomposition processes as guidance. The few-shot prompt for decomposition is provided in Table 7. Only text is used in the few-shot prompt, without images. The decomposer determines the number of sub-questions needed. The few-shot prompt for the second-iteration decomposition is shown in Table 8 For paraphrasing, we use the same samples with manually written paraphrased questions. The few-shot prompt for paraphrasing is provided in Table 9. The remaining datasets are approached with a zero-shot strategy. We use OpenHermes-2.5-Mistral-7B555https://huggingface.co/teknium/OpenHermes-2.5-Mistral-7B as the LLM Agent for reasoning. We evaluate three VLMs: LLaVA1.5-7B Liu et al. (2023), Idefics2-8B Laurençon et al. (2024), and InternVL Chen et al. (2024b), all operating under a zero-shot setting across all datasets. Since all datasets are multiple-choice QA tasks or short answers, we use string matching for answer consistency. For baseline threshold settings:
@@ A.4 Case Study
Method | Sub-Questions/ | Total Samples | Time per Sample |
Paraphrased-Questions per Sample | Used | (Seconds) | |
VCR | |||
[…]
Perplexity of Direct Answer, Generated Confidence | N/a | 1000 | 0.22 |
Self-Consistency based on Paraphrase | 4 | 1000 | 4.80 |
[…]
Question Pre-Decomposition | 3.9 | 1000 | 3.96 |
Sub-Question Answering and VLM Agent Reasoning | 3.9 | 1000 | 0.84 |
LLM Agent Reasoning | 3.9 | 1000 | 0.18 |
Second Question Decomposition | 4.1 | 366 | 4.09 |
Second Sub-Question Answering and VLM Agent Reasoning | 4.1 | 366 | 0.93 |
Second LLM Agent Reasoning | 4.1 | 366 | 0.20 |
[…]
DeCC Expected Value (VLM-Agent Consistency (2 Iterations)) | 8.0 | 1000 | 9.82 |
DeCC Expected Value (LLM-Agent Consistency (2 Iterations)) | 8.0 | 1000 | 10.20 |
AOKVQA | |||
[…]
Perplexity of Direct Answer, Generated Confidence | N/a | 1000 | 0.16 |
Self-Consistency based on Paraphrase | 4 | 1000 | 3.90 |
[…]
Question Pre-Decomposition | 3.2 | 1000 | 3.36 |
Sub-Question Answering and VLM Agent Reasoning | 3.2 | 1000 | 0.54 |
LLM Agent Reasoning | 3.2 | 1000 | 0.10 |
Second Question Decomposition | 3.9 | 253 | 3.79 |
Second Sub-Question Answering and VLM Agent Reasoning | 3.9 | 253 | 0.66 |
Second LLM Agent Reasoning | 3.9 | 253 | 0.12 |
[…]
DeCC Expected Value (VLM-Agent Consistency (2 Iterations)) | 7.1 | 1000 | 8.35 |
DeCC Expected Value (LLM-Agent Consistency (2 Iterations)) | 7.1 | 1000 | 8.57 |
A.5 Computational Costs Analysis
DeCC requires multiple steps – question decomposition, answering the decomposed questions, comparing consistency between the answers. This raises concerns about DeCC being computationally more expensive than other approaches such as perplexity, generated numerical confidence scores, or self-consistency. To address this concern, we provide detailed computational costs for DeCC and compare it with that of baselines. Since the computational cost in DeCC is influenced by the number of sub-questions generated, we conduct this analysis on the VCR and AOKVQA, as these two datasets represent the median number of generated sub-questions among all benchmarks. All statistics are computed using Idefics2-8B Laurençon et al. (2024) as the candidate VLM over a randomly chosen sample of 1,000 instances with a mean of three runs. For LLM Agent reasoning, we use vLLM Kwon et al. (2023) for speeding up inference speed. Others are implemented using the Transformers toolkit666https://huggingface.co/docs/transformers/en/index from Hugging Face without any speeding-up strategies. The consistency comparison stage requires only string comparisons, which is very fast (1 to 2 seconds for 1000 samples), so we do not list the time. We report the computational costs analysis in Table 6. For baselines such as perplexity of direct answer and generated confidence, the time cost is equal to direct answering, i.e., 0.22 seconds per sample for VCR and 0.16 seconds per sample for AOKVQA. For the baseline Self-Consistency based on Paraphrase, the consumed time is similar to the first iteration of DeCC, 4.8 seconds per sample for VCR and 3.9 seconds per sample for AOKVQA. For the first iteration of DeCC, the most time-consuming stage is the Question Pre-Decomposition stage for generating the sub-questions, taking 3.96 seconds per sample for VCR and 3.36 seconds per sample for AOKVQA. However, this process needs to be performed only once per sample before evaluation and can then be reused to evaluate multiple VLMs, thus mitigating the impact on evaluation speed. The Sub-Question Answering and VLM Agent Reasoning + LLM Agent Reasoning takes 1.02 seconds per sample for VCR and 0.64 seconds per sample for AOKVQA. The second iteration of DeCC takes 5.22 seconds per sample for VCR and 4.57 seconds per sample for AOKVQA. However, the second decomposition is required only for samples where there is a disagreement between VLM Agent Consistency and LLM Agent Consistency from the first iteration, rather than for all samples. For example, only 25.3% of samples in AOKVQA and 36.6% of samples in VCR require the second iteration. Thus, the total time for the second iteration is significantly less than the first iteration. Overall the expected total time of DeCC Multi-Agent Setting taking into account both first and second round of decompositions is 6.89 seconds per sample for VCR and 5.16 seconds per sample for AOKVQA. Note that for VLM-Agent Consistency (2 Iterations) and LLM-Agent Consistency (2 Iterations), the expected total time is larger than the multi-agent setting because we need to do the second round of decomposition for all samples. With engineering optimizations and parallel processing, the computing time of DeCC can be further reduced. Note that DeCC does not require any training or tuning of thresholds while other baselines such as perplexity, generated numerical confidence scores, or self-consistency require tuning certain thresholds, necessitating annotated datasets. For example, we used 1,000 samples for threshold tuning in our experiments. Data annotation is usually time-consuming and costly, averaging 10 seconds per sample777CloudResearch: A Simple Formula for Predicting the Time to Complete a Study on Mechanical Turk with further filtering requests, and for specialized domains, this time can increase significantly, thus increasing the hidden costs. Compared with the data annotation, the time cost of DeCC is significantly lower. In conclusion, despite the additional computation introduced by DeCC compared to baselines, our comprehensive experimental evaluation validates the effectiveness and practicality of DeCC across various benchmarks and VLMs.
@@ A.5 Computational Costs Analysis
Paraphrased question 1: Is this statement entailment, neutral, or contradiction based on the image? Statement: ‘A teacher is late to class’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 2: Is this statement entailment, neutral, or contradiction based on the image? Statement: ‘A professor is tardy to class’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 3: Is this statement entailment, neutral, or contradiction based on the image? Statement: ‘A professor is not on time for class’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 4: Is this statement entailment, neutral, or contradiction based on the image? Statement: ‘A teacher is not punctual for class’ Options: A: entailment, B: neutral, C: contradiction. |
@@
Main Question: Is this statement entailment, neutral, or contradiction based on the image? Statement: ‘Two children play in the park.’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 1: Is this statement entailment, neutral, or contradiction based on the image? Statement: ‘Two kids play in the park.’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 2: Is this statement entailment, neutral, or contradiction based on the image? Statement: ‘Two children are playing in the park.’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 3: Is this statement entailment, neutral, or contradiction based on the image? Statement: ‘Two kids are playing in the park.’ Options: A: entailment, B: neutral, C: contradiction. |
Paraphrased question 4: Is this statement entailment, neutral, or contradiction based on the image? Statement: ‘There are two children playing in the park.’ Options: A: entailment, B: neutral, C: contradiction. |